# Can Human Drivers and Connected Autonomous Vehicles Co-exist in Lane-Free Traffic? A Microscopic Simulation Perspective

### v1

@@ 2.4.1 Adaptive Potential Lines (APL) Corridors
The APL controller primarily compresses the PLs into the lateral spaces between HDVs, as shown in Figure 3. This adjustment enables CAVs to overtake slow-moving HDVs and enhances the overall flow of CAVs. Since CAVs already move in a coordinated manner due to the PL controller, they can also overtake HDVs in a coordinated way by simply modifying the assigned PL position ($y_{pl,i}$) around HDVs. Additionally, these areas can create long road regions where CAVs are longitudinally dominated, allowing them higher chances to exert artificial forces on the downstream CAVs and accelerate without being blocked by HDVs. This functions as a sort of corridor for CAVs with comparatively higher freedom and, thus, is referred to as the APL corridor in the study.
To determine the areas where APL corridors can be formed, the method first divides the road into a set of regions $R$ in the longitudinal axis, with $x^{l}_{r}$ and $x^{h}_{r}$ denoting the longitudinal positions of the start and end of the region $r\in R$, respectively. A region starts at a distance $x^{th}_{j}$ behind the HDV $j$ and extends to its front in the longitudinal axis. $x^{th}_{j}$ provides additional safety distance for the CAVs to orient themselves on the modified PLs without excessive deceleration near the HDVs. This region does not extend beyond the front of the HDV to avoid influencing CAVs in front, which is problematic since any HDV approaching from behind would force the CAV to move away from its front, giving unnecessary priority to HDVs. If multiple regions overlap in the longitudinal direction, they are merged to form a single region — an extended APL corridor. Figures 3(b) and 3(c) show examples of such longitudinal overlap and extended APL corridor.
@@ 2.4.2 Activation Condition for APL Corridors
The previous section focused on the potential areas for APL corridors. The two critical factors for applying APL corridors are (1) determining $x^{th}_{j}$ and (2) determining the HDVs whose surroundings should be included for APL corridors. The paper evaluates four approaches that mainly differ in these two aspects, as described below:
@@
    Neighbouring Speed based Constant Margin (NSCM): Similar to CM method, this method uses the same $X_{CM}$ for all HDVs; however, the vicinity (i.e. longitudinal distance $X_{CM}$ behind an HDV) of an HDV is only included into APL corridor if its speed is less than the average speed of the surrounding vehicles. For an HDV $j$, the average surrounding speed, denoted by $v^{s}_{j}$, is computed by taking the mean longitudinal speed of all vehicles (both CAVs and HDVs) whose positions do not laterally overlap with the boundaries of $j$ and are within a longitudinal distance of 20 m behind $j$. The idea here is that the APL should only be applied in the surroundings of an HDV if it is blocking the vehicles behind it, which can be detected if the vehicles with different lateral positions than $j$ are moving faster than $j$.
@@
    Follower-based Adaptive Margin (FAM): The PL controller in this study uses the leader-follower relation along with safe acceleration as a hard constraint to avoid collisions (Eq. 12). The HDV can potentially hinder any CAV that has an HDV leader. FAM uses this characteristic to apply the APL. It first checks if the HDV $j$ is the leader of any CAV within a longitudinal distance of $X_{AM}$ behind the HDV. Then, it checks if the speed of $j$ is lower than the average speed of the surrounding vehicles (the same as in the NSCM strategy). If both conditions are fulfilled, $x^{th}_{j}$ is set as the distance from the back of HDV to the back of the follower CAV. The purpose here is to only apply APL behind an HDV up to a point that allows the overtaking of the follower CAV. Thus, $x^{th}_{j}$ in this strategy is not constant; rather, it depends on the distance of the follower CAV, as reflected in its name.
@@ 3 Experimental Setup
The following values are used for the model parameters. For the human model, $\Delta s$, $\lambda$ and $L_{th}$ are set to 0.1 $m$, 0.1, and 10, respectively. The values of $\Delta s$ and $\lambda$ are based on the values used in Mathew et al. (2015), while $L_{th}$ is not explicitly mentioned by the authors, and thus, the value is chosen by experiments to allow sufficient strip changing maneuvers. For calculating $v^{safe}_{x,i}$, the maximum deceleration $A^{-}_{x}$ is set as -4.5 $m/s^{2}$. $\tau$ is fixed as 0.5 $s$ for all CAVs. However, to bring variety to the individual driving style of each human driver, for HDVs, $\tau$ is drawn from a normal distribution with a mean and standard deviation of 1.5 and 0.5, respectively. The acceleration ability $A^{+}_{x}$ of all vehicles is set to 2.6 $m/s^{2}$. The values of both $A^{-}_{x}$ and $A^{+}_{x}$ are taken from the standard SUMO configuration. $\Delta D_{front}$ and $\Delta D_{back}$ are set to 50 $m$. For the APL controller, $X_{CM}$ and $X_{AM}$ are set to 40 $m$ by default unless explicitly stated otherwise.
@@ 4.1 Impact of HDVs on the LFT performance
The study first analyses the impact of different penetration rates of HDVs on the overall LFT performance, especially in terms of traffic flow and average speed, as shown in Figure 4. The first thing to observe is the significant difference in the road capacities (the highest traffic flow in Figure 4) with all-HDVs and all-CAVs scenarios. With 100% HDVs, the road capacity is limited to only 8,233 $veh/h$ (speed: 23.2 $m/s$) at a density of only 100 $veh/km$, which is roughly equivalent to the flow of lane-based traffic with four lanes (approximately 2000 $veh/h/lane$ Kesting and Treiber (2013)). Notably, the road width used in the simulation (10.2 meters) is typically divided into three lanes in lane-based traffic management. However, since the lanes are usually wider than the vehicles, removing the lanes allowed four vehicles to fit within the same road width, resulting in a flow roughly equivalent to four lanes. Nevertheless, this flow value also depends on the tolerance of individual drivers for changing lateral positions. For instance, in the experiments conducted for this study, the maximum road capacity with all-HDVs decreased from 8,233 $veh/h$ to 6,973 $veh/h$ when the threshold for benefits of changing strip ($L_{th}$) was reduced from 10 to 0.1.
In contrast, the scenario with all-CAVs achieves a significantly higher road capacity of 20,800 $veh/h$ (speed: 23.1 $m/s$) at a density of 250 $veh/km$ — an increase of almost 153% in road capacity over all-HDVs scenarios. The key difference between HDVs and the LFT is the existence of artificial forces, indicating that the coordinated movement of CAVs through artificial forces and virtual PLs plays a crucial role in achieving higher flow. Additionally, despite using the same desired speeds for individual vehicles in both scenarios, the average speed observed in the all-CAVs scenario is significantly higher across all vehicle densities.
Figure 4: Traffic flow rate and mean speed of PL controller with different HDV penetration rates. The 0% and 100% penetration rates mark all-CAVs and all-HDVs scenarios, respectively.
Figure 5: Traffic flow rate and mean lateral speeds for increasing penetration of HDV penetration into LFT.
(a) 250 veh/km
@@
Figure 4 also shows that as the penetration rate of HDVs increases, the flow of LFT is severely affected. Even a small percentage of HDVs can significantly lower LFT performance: with just a 5% penetration rate, road capacity drops to 17,252 $veh/h$ (a drop of nearly 16%), and with a 10% penetration rate, it drops to 14,162 $veh/h$ (a drop of nearly 32%). This phenomenon is further analyzed in Figure 5, illustrating traffic flow and lateral speeds against increasing HDVs penetration rates. It is observed that the performance drop is significantly high for the initial 0-20% HDVs penetration. For individual vehicle densities above 200 $veh/km$, the traffic flow is almost halved at 20% penetration, which continues to drop further at a lower rate until 40% penetration. Beyond this point, the curves are almost flat without much performance gain over all-HDVs scenarios. The performance decline is more pronounced at higher vehicle densities.
The above phenomena can be further understood by looking at the mean lateral speed in Figure 5. First, with all-CAVs scenario, the vehicles can better coordinate their motion and stay on their respective PLs, leading to lower lateral movements. For all-CAVs scenarios, it is also observed that the vehicles generally show higher lateral movements at lower vehicle densities compared to higher vehicle densities. This can be explained by the higher available space at lower densities, causing the overtaking vehicles to fully utilize their artificial forces to nudge the vehicles in the front. As the vehicle densities increase, the nudging possibility decreases due to the artificial forces from the surrounding vehicles countering each other. Nonetheless, the vehicles can still coordinate their movements at higher densities, and flow remains significantly higher than all-HDV scenarios.
@@
Figure 5 also shows that the lateral movements continue to increase with higher penetration rates of HDVs till a saturation point is reached, after which they begin to decline again. As the vehicle densities increase, the saturation point occurs at a lower penetration rate of HDVs. This phenomenon can be explained as follows: at lower penetration rates, the upstream CAVs can better adjust their positions in front of the hindrances caused by HDVs since the CAVs can influence other CAVs, leading to better utilization of space and higher lateral movement; however, as the proportion of HDVs increases, CAVs are unable to do that leading to lower lateral movement and higher gaps between vehicles. This is visible by comparing Figure 7(c), Figure 7(d), where the portions of road with tightly packed CAVs are larger than scenarios with 5% HDVs scenario. This happens till the saturation point, after which the proportion of HDVs is so high that the CAVs find limited opportunity to exert artificial forces, and consequently, the lateral movement and the efficient road space utilization decreases, as illustrated in Figure 7(e). This saturation point happens earlier for higher densities because the number of HDVs needed to cause this lateral movement decline is achieved earlier for higher vehicle densities.
@@
(b) All-CAVs scenario at $t=10$ min.
(c) 5% HDVs at $t=10$ min.
(d) 10% HDVs at $t=10$ min.
(e) 50% HDVs at $t=10$ min.
(f) 100% HDVs at $t=10$ min.
@@
The above phenomena can also be observed from the development of lateral positions of CAVs over time, as shown in Figure 8. The behavior of the PL controller with all-CAVs scenario is discussed first. The PL controller tries to minimize the lateral movement by assigning PL according to desired speeds. At the beginning of the simulation, there are usually higher lateral movements since all CAVs try to reach their assigned PLs. Over time, the CAVs settle down at a lateral position and the lateral movements are minimized subsequently. It is possible that these settling points are not exactly on the assigned PL and are slightly shifted, according to the location of the PL and the artificial forces from surrounding vehicles. It is observed that this deviation of the settling point is larger near the road boundaries, as shown by all-CAVs scenarios in Figure 8. The CAVs with assigned PLs near the boundaries experience artificial forces from the surrounding vehicles from one side only, and in the absence of artificial boundary forces to counter them (due to the modifications of PL controller in the current study), these CAVs are pushed further towards the boundaries. In contrast, the CAVs with assigned PLs near the center, experience artificial forces from both sides, making it easier for them to stay on the assigned PL. However, at higher vehicle densities, even the CAVs at the center may have significant deviation due to the CAVs not finding enough space and opportunity to remain on the assigned PL, as shown by the all-CAVs scenario in Figure 8(b).
(a) 250 $veh/km$

### v2

@@ 3.4.1 Adaptive Potential Lines (APL) Corridors
The APL controller primarily compresses the PLs into the lateral spaces between HDVs, as shown in Figure 4. This adjustment enables CAVs to overtake slow-moving HDVs and enhances the overall flow of CAVs. Since CAVs already move in a coordinated manner due to the PL controller, they can also overtake HDVs in a coordinated way by simply modifying the assigned PL position ($y_{pl,i}$) around HDVs. Additionally, these areas can create long road regions where CAVs are longitudinally dominated, allowing them higher chances to exert artificial forces on the downstream CAVs and accelerate without being blocked by HDVs. This functions as a sort of corridor for CAVs with comparatively higher freedom and, thus, is referred to as the APL corridor in the study.This study assumes that APL calculations are performed centrally by dedicated controllers positioned along the roadside. These controllers estimate the locations and speeds of HDVs and transmit the relevant APL information to CAVs. Alternatively, a decentralized approach could be employed, where individual CAVs contribute to the APL calculations. In this scenario, CAVs would share their location and speed estimates with one another via V2V communication, allowing a distributed algorithm to fuse the data for improved accuracy in HDV estimation. The fused data would then be used to perform the APL calculations and disseminate the resulting information to nearby CAVs.
To determine the areas where APL corridors can be formed, the method first divides the road into a set of regions $R$ in the longitudinal axis, with $x^{l}_{r}$ and $x^{h}_{r}$ denoting the longitudinal positions of the start and end of the region $r\in R$, respectively. A region starts at a distance $x^{th}_{j}$ behind the HDV $j$ and extends to its front in the longitudinal axis. $x^{th}_{j}$ provides additional safety distance for the CAVs to orient themselves on the modified PLs without excessive deceleration near the HDVs. This region does not extend beyond the front of the HDV to avoid influencing CAVs in front, which is problematic since any HDV approaching from behind would force the CAV to move away from its front, giving unnecessary priority to HDVs. If multiple regions overlap in the longitudinal direction, they are merged to form a single region — an extended APL corridor. Figures 4(b) and 4(c) show examples of such longitudinal overlap and extended APL corridor.
@@ 3.4.2 Activation Condition for APL Corridors
The previous section focused on the potential areas for APL corridors. The two critical factors for applying APL corridors are (1) determining $x^{th}_{j}$ and (2) determining the conditions needed for an HDV to include its surroundings into APL corridors. The paper evaluates four approaches that mainly differ in these two aspects. Figure 5 shows the parameters used by these methods. The APL methods used are described below:
Figure 5: Schematic to show parameters used for APL methods.
Method The value of $x_{j}^{th}$ Condition for Corridor Formation
CM Fixed as $X_{CM}$ Unconditional
NSCM Fixed as $X_{CM}$ HDV’s speed is less than surrounding speed, i.e. $v_{x,j}<v_{j}^{s}$.
FAM Dependent on the distance of the CAV following HDV, i.e., $\Delta x_{f}$
(1) $v_{x,j}<v_{j}^{s}$
(2) The HDV leads a CAV within $X_{AM}$ distance.
SVAM Dependent on the distance of the CAV following HDV, i.e. $\Delta x_{f}$
(1) $v_{x,j}<v_{j}^{s}$
(2) The HDV leads a CAV within $X_{AM}$ distance.
(3) The follower CAV’s velocity restricted by the
safe velocity, i.e., $v_{x,i}<1.05v_{x,i}^{safe}$
Table 2: Summary of the APL methods.
@@
    Neighbouring Speed based Constant Margin (NSCM): Similar to CM method, this method uses the same $X_{CM}$ for all HDVs; however, the vicinity (i.e. longitudinal distance $X_{CM}$ behind an HDV) of an HDV is only included into APL corridor if its speed is less than the average speed of the surrounding vehicles. For an HDV $j$, the average surrounding speed, denoted by $v^{s}_{j}$, is computed by taking the mean longitudinal speed of all vehicles (both CAVs and HDVs) whose positions do not laterally overlap with the boundaries of $j$ and are within a longitudinal distance of 20 m behind $j$. Figure 5 provides an example of calculating $v^{s}_{j}$. The idea here is that the APL should only be applied in the surroundings of an HDV if it is blocking the vehicles behind it, which can be detected if the vehicles with different lateral positions than $j$ are moving faster than $j$, i.e., $v_{x,j}<v_{j}^{s}$.
@@
    Follower-based Adaptive Margin (FAM): The PL controller in this study uses the leader-follower relation along with safe acceleration as a hard constraint to avoid collisions (Eq. 12). The HDV can potentially hinder any CAV that has an HDV leader. FAM uses this characteristic to apply the APL. It first checks if the HDV $j$ is the leader of any CAV within a longitudinal distance of $X_{AM}$ behind the HDV. Then, it checks if the speed of $j$ is lower than the average speed of the surrounding vehicles (the same as in the NSCM strategy). If both conditions are fulfilled, $x^{th}_{j}$ is set as the distance from the back of HDV to the back of the follower CAV, shown as the distance $\Delta x_{f}$ in Figure 5. The purpose here is to only apply APL behind an HDV up to a point that allows the overtaking of the follower CAV. Thus, $x^{th}_{j}$ in this strategy is not constant; rather, it depends on the distance of the follower CAV, as reflected in its name.
@@
Table 2 summarizes the parameters and conditions used by each APL method.
@@ 4 Experimental Setup
The following values are used for the model parameters. For the human model, $\Delta s$, $\lambda$ and $L_{th}$ are set to 0.05 $m$, 0.1, and 10, respectively. The values of $\Delta s$ and $\lambda$ are based on the values used in Mathew et al. (2015), while $L_{th}$ is not explicitly mentioned by the authors, and thus, the value is chosen by experiments to allow sufficient strip changing maneuvers. For calculating $v^{safe}_{x,i}$, the deceleration $A^{-}_{x}$ is set as -1.5 $m/s^{2}$ to ensure smooth deceleration and minimize jerk. To account for situations requiring more aggressive braking of HDVs due to changes in lateral positions, the higher deceleration threshold $A^{-}_{critical}$ is set as -2.6 $m/s^{2}$. Note that $A^{-}_{critical}$ is only used by HDV as the LFT forces already cater for it for CAVs. $\tau$ is fixed as 0.5 $s$ for all CAVs. However, to bring variety to the individual driving style of each human driver, for HDVs, $\tau$ is drawn from a normal distribution with a mean and standard deviation of 1.5 and 0.5, respectively. The minimum safety gap $g_{o}$ is set to 2 m. The desired acceleration $A^{+}_{x}$ of all vehicles is also set to 1.5 $m/s^{2}$. For CAVs, the acceleration limits $(A_{x}^{min},A_{x}^{max})$ are set to (-4.5 $m/s^{2}$, 2.6 $m/s^{2}$) and $(A_{y}^{min},A_{y}^{max})$ are set to (-1.5 $m/s^{2}$, 1.5 $m/s^{2}$). The jerk limits $(J_{x}^{min},J_{x}^{max})$ and $(J_{y}^{min},J_{y}^{max})$ are set to (-2.0 $m/s^{3}$, 2.0 $m/s^{3}$). The acceleration and jerk values are mainly derived from Bae et al. (2019) for comfortable and safe driving experiences. $\Delta D_{front}$ and $\Delta D_{back}$ are set to 100 $m$. The weights for nudging and repulsive forces, i.e., $w_{n}$ and $w_{r}$, are set as 1.0 and 0.5, while the controller gains for the PL controller $K_{pl}$ and $K_{pl,v}$ are set as 0.02 and 0.65, respectively. For boundary control, the values of $K_{b1}$ and $K_{b2}$ are set as 4 and 3.75, respectively. $X_{CM}$ and $X_{AM}$ are set to 40 $m$ by default unless explicitly stated otherwise.
@@ 5.1 Impact of HDVs on the LFT performance
The study first analyses the impact of different penetration rates of HDVs on the overall LFT performance, especially in terms of traffic flow and average speed, as shown in Figure 6. The first thing to observe is the significant difference in the road capacities (the highest traffic flow in Figure 6) with all-HDVs and all-CAVs scenarios. With 100% HDVs, the road capacity is limited to only 8,100 $veh/h$ (speed: 22.5 $m/s$) at a density of only 100 $veh/km$, which is roughly equivalent to the flow of lane-based traffic with four lanes (approximately 2000 $veh/h/lane$ Kesting and Treiber (2013)). Notably, the road width used in the simulation (10.2 meters) is typically divided into three lanes in lane-based traffic management. However, since the lanes are usually wider than the vehicles, removing the lanes allowed four vehicles to fit within the same road width, resulting in a flow roughly equivalent to four lanes. Nevertheless, this flow value also depends on the tolerance of individual drivers for changing lateral positions. For instance, in the experiments conducted for this study, the maximum road capacity with all-HDVs decreased from 8,100 $veh/h$ to 7,400 $veh/h$ when the threshold for benefits of changing strip ($L_{th}$) was reduced from 10 to 0.1.
In contrast, the scenario with all-CAVs achieves a significantly higher road capacity of 16,700 veh/hveh/h [^1] The 16,700 $veh/h$ road capacity is observed using the baseline parameter values mentioned in the paper. A higher flow is still possible with further tuning of the parameters. For example, by setting $w_{n}$ and $w_{r}$ to 1.5 and 1.0, respectively, a road capacity of 18,000 $veh/h$ was observed. However, this led to significantly higher CAV movements under mixed scenarios with HDVs. The road capacity increases even further by adjusting the minimum safety gap $g_{o}$ and reaction time $\tau$. For example, it increases to almost 20,000 $veh/h$ by setting $g_{o}$ to 0, of course, at the cost of significantly higher safety risks. (speed: 23.1 $m/s$) at a density of 200 $veh/km$ — an increase of almost 106% in road capacity over all-HDVs scenarios. The key difference between HDVs and the LFT is the existence of artificial forces, indicating that the coordinated movement of CAVs through artificial forces and virtual PLs plays a crucial role in achieving higher flow. Additionally, despite using the same desired speeds for individual vehicles in both scenarios, the average speed observed in the all-CAVs scenario is significantly higher across all vehicle densities.
It is also worthwhile to compare the obtained maximum capacity with conventional lane-based traffic. For a 10.2 m three-lane road, conventional lane-based traffic reaches about 6,000 $veh/h$, whereas lane-based CAVs with Cooperative Adaptive Cruise Control (CACC) could theoretically achieve up to 12,000 $veh/h$ (assuming 3500–4000 $veh/h/lane$) under ideal conditions Papamichail et al. (2019); Shladover et al. (2012); Vander Werf et al. (2002). This remains well below the 16,700 $veh/h$ observed for all-CAV LFT. In practice, lane-based CACC capacity would be even lower once lane-changing disruptions are considered, whereas LFT avoids these losses through continuous lateral positioning.
Figure 6: Traffic flow rate and mean speed of PL controller with different HDV penetration rates. The 0% and 100% penetration rates mark all-CAVs and all-HDVs scenarios, respectively. The shaded area shows the standard deviation.
Figure 7: Traffic flow rate and mean lateral speeds for increasing penetration of HDV into LFT.
(a) 200 veh/km
@@
Figure 6 also shows that an extremely small HDV penetration of 1% (i.e., 1-4 HDV depending on vehicle density) has a comparatively limited effect on the LFT flow. This HDV penetration is relevant to scenarios where some exceptional vehicles, such as emergency vehicles or towed cars, have to be manually driven. However, with even slight increase in HDV penetration, the LFT flow is severely affected: with just a 5% penetration rate, road capacity drops to 13,300 $veh/h$ (a drop of nearly 20%), and with a 10% penetration rate and the same vehicle density of 200 $veh/h$, the flow is decreased to 11,700 $veh/h$ (a drop of nearly 30%). Interestingly, at 10% HDVs the maximum capacity remains nearly the same as with 5% HDVs, but the critical density decreases to 150 $veh/km$. This shift occurs because higher HDV penetration reduces average traffic speed, limiting CAVs’ ability to overtake or influence HDV behavior. Nevertheless, the PL controller still maintains similar capacity at 10% HDVs as at 5%, although at the lower critical density. This demonstrates that HDV penetration not only degrades LFT performance but can also shift the fundamental diagram, depending on CAV–HDV interaction. This phenomenon is further analyzed in Figure 7, illustrating traffic flow against increasing HDVs penetration rates. It is found that the above drastic traffic flow drop with 5% and 10% HDVs penetration is consistent for all vehicle densities above $150\penalty 10000\ veh/km$. In fact, the performance drop is significantly high for the initial 0-20%, where at 20% penetration rate, the flow already drops by 40%. Beyond this point, the curves are almost flat with limited performance gain over all-HDVs scenarios: from 40-100% the performance drops further by only 10%, reaching a drop of nearly 60% for all-HDVs scenario.
The above phenomena can be better understood by examining vehicle movements along the lateral axis. In general, when slower vehicles do not obstruct faster ones, vehicles with higher desired speeds maintain their trajectory with minimal lateral deviation, resulting in smoother traffic flow. Conversely, when slower vehicles impede faster vehicles, the latter must frequently adjust their lateral position to overtake, leading to increased lateral movement. We quantify this behavior using the mean of the absolute values of lateral speeds, referred to as the mean lateral speed in this study. As shown in Figure 7, in the all-CAVs scenario, vehicles are able to better coordinate their movements and remain on their respective PLs, resulting in reduced lateral movement and a correspondingly lower mean lateral speed. For all-CAVs scenarios, it is also observed that the vehicles generally show higher lateral movements at lower vehicle densities compared to higher vehicle densities. This can be explained by the higher available space at lower densities, causing the overtaking vehicles to fully utilize their artificial forces to nudge the vehicles in the front. As the vehicle densities increase, the nudging possibility decreases due to the artificial forces from the surrounding vehicles countering each other. Nonetheless, the vehicles can still coordinate their movements at higher densities, and flow remains significantly higher than all-HDV scenarios.
@@
Figure 7 also shows that the lateral movements continue to increase with higher penetration rates of HDVs till a saturation point is reached, after which they begin to decline again. As the vehicle densities increase, the saturation point occurs at a lower penetration rate of HDVs. This phenomenon can be explained as follows: at lower penetration rates, the upstream CAVs can better adjust their positions in front of the hindrances caused by HDVs since the CAVs can influence other CAVs, leading to better utilization of space and higher lateral movement; however, as the proportion of HDVs increases, CAVs are unable to do that, leading to lower lateral movement and higher gaps between vehicles. This is visible by comparing Figure 9(c), Figure 9(d), where the portions of road with tightly packed CAVs are larger than scenarios with 5% HDVs scenario. This happens till the saturation point, after which the proportion of HDVs is so high that the CAVs find limited opportunity to exert artificial forces, and consequently, the lateral movement and the efficient road space utilization decreases, as illustrated in Figure 9(f). This saturation point happens earlier for higher densities because the number of HDVs needed to cause this lateral movement decline is achieved earlier for higher vehicle densities.
@@
(b) All-CAVs scenario at $t=50$
min.
(c) 5% HDVs at $t=50$ min.
(d) 10% HDVs at $t=50$ min.
(e) 20% HDVs at $t=50$ min.
(f) 50% HDVs at $t=50$ min.
(g) 100% HDVs at $t=50$ min.
@@
The above phenomena can also be observed from the development of lateral positions of CAVs over time, as shown in Figure 10. The behavior of the PL controller with all-CAVs scenario is discussed first. The PL controller tries to minimize the lateral movement by assigning PL according to desired speeds. At the beginning of the simulation, there are usually higher lateral movements since all CAVs try to reach their assigned PLs. Over time, the CAVs settle down at a lateral position and the lateral movements are minimized subsequently. It is possible that these settling points are not exactly on the assigned PL and are slightly shifted, according to the location of the PL and the artificial forces from surrounding vehicles. It is observed that this deviation of the settling point is larger near the road boundaries, as shown by all-CAVs scenarios in Figure 10. The CAVs with assigned PLs near the boundaries experience artificial forces from the surrounding vehicles from one side only, and in the absence of any artificial forces to counter them, these CAVs are pushed further towards the boundaries. In contrast, the CAVs with assigned PLs near the center, experience artificial forces from both sides, making it easier for them to stay on the assigned PL. However, at higher vehicle densities, even the CAVs at the center may have significant deviation due to the CAVs not finding enough space and opportunity to remain on the assigned PL, as shown by the all-CAVs scenario in Figure 10(b).
(a) 200 $veh/km$